OpenStudy (cutiecomittee123):

Help me solve this equation -3=log"b" (1/216) the first step I took was this: b^-3-(1/216)

OpenStudy (anonymous):

sorry this far to much for me

OpenStudy (cutiecomittee123):

Haha okay thanks anyways

Nnesha (nnesha):

b is a base right ? and is that b^-3 minus or equal 1/216 ?

OpenStudy (cutiecomittee123):

Yes b is a base. b^-3 = 1/216

OpenStudy (cutiecomittee123):

@Nnesha

Nnesha (nnesha):

okay you already know the exponent rule negative exponent so ....?

OpenStudy (cutiecomittee123):

so with negative exponents you do 1/Xm

Nnesha (nnesha):

i guess u meant to say 1/x^m m exponent right so yes

OpenStudy (cutiecomittee123):

Yea I did

OpenStudy (cutiecomittee123):

Okay so I wrote it as 1/b^3 = 1/216

OpenStudy (cutiecomittee123):

is that right?

Nnesha (nnesha):

yes that's right

Nnesha (nnesha):

now cross multiply

OpenStudy (cutiecomittee123):

can you show me how to do that, I am having a brain fart

Nnesha (nnesha):